Maria bicycles 7km/h faster than Javier. In the same time it takes Javier to bicycle 51km Maria can bicycle 72km. How fast does each bicyclist travel?
also
Find the LCM of (4+3y), (16-9y^2), and (4-3y)

Answer by mananth(16949)

Javier B x mph
Maria A 7 faster

Javier 51 miles
Maria 72 miles
Javier time = Maria time

51 / x = 72 /( x + 7 )
cross multiply
51 ( x + 7 ) = 72 x
51 x + 357 = 72 x
72 x - 51 x = 357
21 x = 357
/ 21
x = 17 mph Javier

24 mph Maria
Find the LCM of (4+3y), (16-9y^2), and (4-3y)
(4+3y), (4+3y)(4-3y), and (4-3y)
LCM = (4+3y)(4-3y)